OpenStudy (anonymous):

How do you differentiate (x+1)/x?

OpenStudy (anonymous):

\[\frac{ x+1 }{ x }\]

OpenStudy (anonymous):

2 ways you can do this: 1) Use the quotient rule 2) Even easier is realizing: (x + 1)/x = 1 + 1/x You can take the derivative of 1, which is 0 and add that 0 (or disregard it) to the derivative of 1/x

OpenStudy (anonymous):

1/x = x^(-1) derivative of x^a = ax^(a-1)

OpenStudy (anonymous):

So, your "a" is -1, so now just substitute.

OpenStudy (anonymous):

Is this making sense to you?